MainActor

a modular animation package Version 1.16 20 July 1993

by Markus Moenig

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1993 Markus Moenig Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies. Permission is granted to copy and distribute modified versions of this manual under the conditions for verbatim copying, provided that the entire resulting derived work is distributed under the terms of a permission notice identical to this one.

1 Warm Up

MainActor is shareware. The program may be freely distributed and copied as long as the following conditions are acknowledged: − The sales price must not be higher than the cost of an (empty) disk plus

a nominal copying fee plus costs for shipping. The total price must not be higher than 6 US$ or 10 DM.

− All parts of the program and the documentation must be complete. The

distribution of single parts is not allowed. − MainActor or parts of it must not be sold in combination with commer-

cial software without the written permission of the author. − Program and documentation must not be changed in any ways. The

repacking of this distribution with different packers is however allowed. − Permission is granted to include this archive in Public-Domain collec-

tions, especially in Fred Fishs Amiga Disk Library (including CD ROM versions of it).

− The author is not responsible for misuse, or damage caused by Main-Actor.

2 Installation

Just copy the MainActor drawer to a device of your choice and add an assign to your user-startup, like

   assign MainActor: <Path>

The default configuration file will clone your Workbench, the positions of the windows are however saved according to my setup and may look strange on yours.

If you have an older version of MainActor installed, make sure to delete it before installing the new version (except perhaps the preferences files).

3 Introduction

3.1 What is MainActor ?

MainActor is a modular animation package for the Amiga. There exist four different kinds of modules:

1) Animation Loaders

2) Animation Savers

3) Picture Loaders

4) Picture Savers this can change in the future.

You can handle up to five projects; every project can have an animation or picture loader, as well as an animation or picture saver. This makes it possible to convert everything in any ways, the number of available modules can be easily upgraded.

3.2 Features

− Intelligent caching/tracking, MainActor can cache your animations and picture lists, if you have not enough memory it will directly access your data from any device. This makes it for example possible to create/edit/play a 400MB animation on your 2MB Amiga.

− If you run OS 3.0 or higher, you will get an extra speed bonus on animations through the use of the new graphics routines. For OS2.0/2.1 users there are specific routines in the modules, which will give them the best playback speed possible on their system (and my knowledge :)).

− Under OS 3.0 you will get the extra features of showing pictures or playing animations in a scrollable, resizable window on your MainActor or Workbench screen. The colors will be properly adjusted to your screen attributes.

− The modules have play routines of their own, so it will be possible to redirect the input/output of special modules to graphic cards or other hardware.

− The play routines support a timecode per frame option. A Sound Mod-

ule (per frame) option will be there in one of the next releases. − MainActor has a totally user reconfigureable GUI. The settings of your

sessions can be saved, this includes the size/position of windows as well as the status of the projects.

− MainActor has an arexx port, nearly all functions can be accessed through it. You can for example scale or dither whole animations on the fly through the use of an image processor, scripts are included.

− MainActor lets you setup and save nearly everything concerning the display mode for playing an animation or showing a picture. For example you can set the X/Y offsets, the display mode ID of your screen (view) and so on.

− MainActor supports localization.

3.3 New features in v1.1

− You can now map your arexx scripts to the function keys. − New IFF-Anim8 32, IFF-Anim8 16, PCX modules. − The new Universal modules allow you to easily load all of the supported

graphic formats. The Universal Picasso-II modules remap the output of all modules directly to the Picasso-II graphics board from Village Tronic.

− MainActor has now an powerful external player named ’MainView’, it can view all animations/pictures for which MainActor has the proper loader module. It is executeable from the wb or cli and supports, per tooltypes or cli arguments, all of MainActors view options.

− MainActor now (optionally) saves icons for your newly created animations or pictures, enabling you to view the animations/pictures by selecting their icons. These icons support the same tooltypes as MainView.

− You can now control how many times MainActor/MainView play ani-

mations. − The new ’Flash’ option allows you to profile your animations.

4 The Menus

You are able to access most of MainActor’s functions through the various menu items and their equivalent shortcuts. Some functions can only be accessed through the menus.

4.1 Project Menu

Open New Project

will open a new project window. You can open up to five project windows, see Section 5.1 [Project Window], page 8.

The attributes of the project window will be set, according to the defaults selectable in the Section 5.5 [MainActor Preferences Window], page 17.

Project

Consists of several submenu items, their functions are identical to the gadgets of the Section 5.2 [Project Control Panel Window], page 10.

Open Prefs

Opens an asl-requester which lets you choose a MainActor preferences file, saved with one of the two following menu items. When you have choosen a prefs file, the display AND all projects and their contents will be closed, MainActor will then be reopened with the settings of the preferences file.

You can also activate MainActor by simply clicking on the icons of the preferences files.

Save Prefs As ...

Saves your current setup of MainActor; nearly everything will be saved : the projects informations( All view settings, the actual project name, the size and position of all windows, the settings of the system preferences ...).

An self-execuing icon will be created for the preferences file. You should save every setup you use frequently !

Save Prefs As Defaults

This will save your setup as the default preferences file, which will be automatically loaded when you start MainActor. The default preferences file is MainActor:Prefs/Default_Prefs. Replace the default by your most used setup.

Iconify

MainActor will be terminated and the MainActor icon will be placed on your workbench. If you want to reactivate MainActor, just click on the icon.

About

Gives you some informations of the current version of MainActor and its author.

Quit

Will close all projects and quit MainActor.

4.2 Text Buffer Menu

MainActor uses the Section 5.6 [Text Buffer Window], page 17 for presenting feedback to the user.

Print Buffer

Prints the informations stored in the Text Buffer Window.

Save Buffer As ...

Will let you choose a file into which the contents of the Text Buffer Window will be saved.

Clear Buffer

Will clear the Text Buffer Window and its buffer.

4.3 Miscellaneous Menu

The menu items of this menu will only be accessable when a project is loaded in the source project. The Set Timecode item will only be activated if an animation project is loaded in the source project.

Select All

Will select all items frames/pictures in your source project.

Select Range

A requester will pop up, just enter the start/stop frame for the range of items to be selected.

DeSelect All

This item will deselect all frames/pictures in your source project.

DeSelect Range

A requester will pop up, just enter the start/stop frame for the range of frames/pictures to be deselected.

Set Timecode

Enter the timecode in the requester which you want to be applied to all selected frames in the source project. Remember to hit return after you typed a new value into the integer gadget. The new timecode will be set immediately. All timecodes are entered in 1/60 of a second, MainActor calculates the right timecode for your vertical beam frequency itself.

4.4 Arexx Menu

See Chapter 7 [The Arexx Interface], page 22,for more informations of MainActors arexx capabilities.

Install Arexx Scripts

This menu item will open a window through which you can map your most used arexx scripts to the function keys. The cycle gadget in the upper part of the window will let you choose the function key on which you want to install an arexx script. The function keys are displayed as F1 ... F20, F10 ... F20 mean the function key + one of the shift keys. You can choose the function key by clicking on the gadget or (the faster method) by simply pressing the appropriate function key. (Remember that you have to press F1 together with the shift key to get the F11 key).

Below the cycle gadget are the Select Script ... gadget and a string gadget. The arexx script which is currently associated to the actual function key displayed in the cycle gadget is visible in the string gadget. You can enter a new arexx script or edit the name of the existing one through the string gadget. You can use the Select Script ... gadget to choose the arexx script via an asl requester.

The three gadgets in the lower part of the window are labeled Save, Use

and Cancel.

Save saves your definitions of the function keys, they will be automatically reloaded the next time you start MainActor.

Use will keep the definitions, but will not save them

Cancel ignores the changes you made to the definitions.

Start Arexx Script

Will let you choose an arexx script, which will then be executed.

4.5 Windows Menu

This menu will let you open/close the different MainActor windows by simply selecting the desired menu item.

5 The Windows

The windows are MainActors main interface to the user. Their use is however optional, you can close all of them (except the Project Windows) and only use the menus. The choice is yours. The position and status of the windows will be saved in the preferences files.

The following is a detailed description of MainActors windows and their gadgets.

5.1 Project Window

A new Project Window will be opend by selecting the Open New Project menu item in the Section 4.1 [Project Menu], page 5.

You can open up to five Project Windows, the number of your Project Window will be displayed in the title bar, in front of the project name, like 1: HAM8_Animation. If no project is loaded, the title is Unused Project.

If you have loaded a project, you can find further informations about it in the screens title bar (when the Project Window is activated). (CPF stands for Color per Frame, CPA for Color per Animation).

The Project Windows are mainly controlled by the Section 5.2 [Project Control Panel Window], page 10.

Project Mode Gadget

Depending on your projects mode, it will be labeled Source Project, Destination Project or DeActivated Project. The Source Project

You can only have one source project. This is the project on which all actions from the Section 5.2 [Project Control Panel Window], page 10 will be executed. The gadgets of the Project Control Panel and the Project View Settings will be setup according to the source project.

The Destination Project This is the project into which all selected data from the source project will be appended through the "Append" gadget of the Project View Settings Window. There can only be one Destination Project.

The DeActivated Projects These projects are waiting to be switched into the source or destination mode. There can be more than one DeActivated Project.

You can change the project mode by clicking on the Project Mode Gadget, for example if you click on the source projects Project Mode Gadget, it will be changed into the destination project and the former destination project will changed into the source project.

If you click on a DeActivated projects Project Mode Gadget, it will change into the source project.

If only one Project Window is open, you cannot change the Project Mode

Gadget.

Everytime you switch on the source project, the gadgets in the Section 5.2 [Project Control Panel Window], page 10 and the Section 5.3 [Project View Settings Window], page 13 will be adjusted to the settings of the new source project.

Caching

This is a simple toggle gadget. If it is checked, it means that your Project is cached or is going to be cached in ram.

If you have not enough ram, you will have to disable it. That means that for example animations will be directly played from your harddisk, which will be much slower than playing them directly from ram.

You can only toggle the status of the Caching gadget when the project window is unused. As soon as you load a project, it will become ghosted.

Select Load Module

When clicking on this Button Gadget, a requester will appear in which you can select the loader module for this project.

The cycling gadget in the requester will let you switch between Picture Loaders or Animation Loaders.

When hitting the Load gadget in the Section 5.2 [Project Control Panel Window], page 10, MainActor executes the animation/picture loader you have selected for the source project.

The name of the current loader module is visible in a text gadget directly under the Select Load Module gadget. Next to the name is a small recessed box with an A or P in it. They stand for Animation or Picture loader.

The Select Load Module gadget will be ghosted as soon as you load a project.

Select Save Module

When clicking on this button gadget, a requester will appear in which you can select the saver module.

The cycling gadget in the requester will let you switch between Picture Savers or Animation Savers.

When hitting the Save gadget in the Section 5.2 [Project Control Panel Window], page 10, MainActor executes the animation/picture saver you have selected for the source project.

The name of the current saver module is visible in a text gadget directly under the Select Save Module gadget. Next to the name is a small recessed box with an A or P in it. They stand for Animation or Picture saver.

The Select Save Module gadget will only be accessable when a project is loaded.

The Scrolling Frames/Picture List

The scroll list will let you choose the frames or pictures you want to execute a certain operation on.

If you have an animation loader for this project window, the scroll list will show you the numbered Frames of the Animation along with their timecodes.

If you have a picture Loader, the scroll list will show you the names of the pictures you have loaded.

Just (de)select the frame/picture by clicking on its text. For multiple selection do not release the mouse button and move over the items to be selected. The list will scroll if you come to the top or bottom border.

See Section 5.5 [MainActor Preferences Window], page 17, you can select how many items shall be displayed in the scroll list. You can also select a special font for it. (The default is Topaz 8 for the font and 8 for the number of items which shall be visible).

5.2 Project Control Panel Window

This window executes its functions on the current source Section 5.1 [Project Window], page 8.

Load

If you have an animation loader selected for the source project, an requester will appear in which you have to select an animation of the appropiate type.

If you have an picture loader selected, the first requester will ask you about the first Picture, after selecting this picture the second qequester will ask you about the last picture. MainActor will now alphabetically sort the pictures and load them. For example, if you have a picture list:

    Amiga_Pic.00001
    Amiga_Pic.00002
    .
    .
    Amiga_Pic.00155

and select Amiga Pic.00055 in the first requester and Amiga Pic.00100 in the next, all pictures from Amiga Pic.00055 to Amiga Pic.00100 will be loaded. The second requester is optional, that means you can abort it by clicking on Cancel if you just want to load one picture.

The Load gadget is only available if you have not loaded a project into the source project yet.

Unload

Is the gadget directly under the Load gadget. It will unload the current project of the source project.

It is only available when you have loaded a project in the source project.

Play

Is next to the Load gadget. It is only available when you have loaded an animation project in the source project.

If you play your animation on a screen, there are two routines available. MainActor checks automatically if you run an OS Version lower than OS 3.0 and will play it the old fashioned way. If you have OS 3.0 or higher you will profit from the new double buffering routines.

The animation will be played according to the settings of the Section 5.3 [Project View Settings Window], page 13.

The following keys can be used to control the playing of the animation:

ESCAPE OR A MOUSEBUTTON

will abort playing the animation

SPACE will stop playing the animation. While the animation is stopped you can use the TAB key to flip to the next frame.

RETURN will enable the timecodes you see in the scroll list of the source

project and which you can change (These timecodes are default).

THE CURSOR KEYS

scroll the display, the offset changes will be shown in the Project View Settings Windows x and y offset gadgets.

N

M If you have activated the Frame Info option, you can cycle the

color of the info text using N and M up and down.

THE FUNCTION KEYS

are used to set the frame rate, they override the defaults. You can reactivate the defaults by pressing RETURN.

F1 sets the timecode to 1/60 Second (60 fps)

F2 sets the timecode to 2/60 Second (30 fps)

F3 sets the timecode to 3/60 Second (20 fps)

F4 sets the timecode to 4/60 Second (15 fps)

F5 sets the timecode to 5/60 Second (12 fps)

F6 sets the timecode to 10/60 Second (6 fps)

F7 sets the timecode to 20/60 Second (3 fps)

F8 sets the timecode to 40/60 Second (1.5 fps)

F9 sets the timecode to 60/60 Second (1 fps)

F10 sets the timecode to 120/60 Second (0.333 fps)

Show

Is the gadget under Play. It will display all the frames/pictures you have selected in the source projects scroll list. This gadget is available for both animation and picture projects.

The frames/pictures will be shown according to the settings of the Section 5.3 [Project View Settings Window], page 13.

The following keys can be used to control the showing of the pictures:

ESCAPE OR A MOUSEBUTTON

will abort showing the pictures.

SPACE will show the next pictures you have selected.

THE CURSOR KEYS

scroll the display, the offset changes will be shown in the Project View Settings Windows x and y offset gadgets.

Save WILL NOT BE USEABLE WITHOUT A KEYFILE

If you have an animation saver activated in your source project, it will ask you about the name of the new animation. Then it will build up the new animation in the appropriate animation Format. It will only save the frames you have SELECTED in the source projects scroll list. If you just want to convert or resave the animation, you have to select all the frames in the scroll list.

If you have a picture saver activated in your source project, a requester will appear in which you have to select the directory and the base name of the pictures.

Example: You have activated an animation loader and a picture saver and loaded an animation. The animation is called Example_Anim. Now you select the first ten frames of the animation and click on the Save gadget. In the requester you type Ram: for the Directory and Example_Pics for the base name. The ten pictures will then be saved as

    Ram:Example_Pics.00001
    Ram:Example_Pics.00002
    .
    .
    Ram:Example_Pics.00010

This is also valid if you have a picture loader and a picture saver.

If you have a picture loader and an animation saver, some more gadgets will appear and ask you for some specific attributes of the new animation:

Palette per Picture and Looping Animation. If you have an animation

loader and saver the saver module will only ask, if you want to have a looping animation, the color per frame setting will be set according to the source animation. (You could change this by executing the Save function through an arexx script). The following combinations are possible:

1. Animation Loader and Animation Saver

2. Animation Loader and Picture Saver

3. Picture Loader and Animation Saver

4. Picture Loader and Picture Saver

In this way, you can easily convert any picture/animation format into any picture/animation format. Append WILL NOT BE USEABLE WITHOUT A KEYFILE

You can append the frames/pictures you have selected in the source project to the destination project. The destination project has to be an animation project.

The source projects animation saver has to belong to the destinations loader module. (For Example the IFF-Anim5 loader and saver). The destination project must NOT be cached.

Some animation types have to fulfill some other attributes, for example the Section 6.1.3 [IFF-Anim5], page 18 format can only append to nonlooping animations. When you want to append frames to a loop Anim5 animation, just resave it and select NO in the loop animation requester, then append the frames.

If you want to add a loop to a non-looping animation, just append two pictures which are identical to the first two, then reload the animation.

See Chapter 6 [The Modules], page 18, for more information about the specific attributes of the modules.

5.3 Project View Settings Window

This window controls the output of the Play and Show gadgets of the Section 5.3 [Project View Settings Window], page 13. The values will be set individually for each project and will be saved with the preferences files.

Frame Info

This switch (when activated) will give you informations about the current frame while playing animations, it is only availaible if the source project contains an animation.

If playing on a screen you will see the info on the top right edge of the screen. When you see no info text, use the M and N keys while playing to cycle the color of the info text up and down.

When playing in a window, the windows title text will display the infos for the current frame.

Mode

This cycle gadget will let you choose between the screen and window mode. In the screen mode the animation/pictures will be shown on a screen (view) of its own.

In the window mode the animation/pictures will be shown in a resizeable, scrollable window on your MainActor screen. THIS OPTION IS ONLY AVAILABLE ON SYSTEMS RUNNING OS3.0+.

The more colors the MainActor screen has available (see Section 5.4 [System Preferences Window], page 15), the better is the dithering of your frame/picture. If you show a 256 color picture on a 4 Color screen, you will be disappointed by the result. The timecodes of the animation will be ignored, the animation will be displayed as fast as possible.

Select Display

You can choose the resolution/view mode of your animation/picture with this gadget (makes only sense in the screen mode). This will override the old value, if you resave the animation/ picture, the new value will be saved.

X/Y Offsets

With these two integer gadgets, you can select the offsets for showing animations/pictures (again, this makes only sense in the screen mode). If you scroll a frame (while playing an animation or showing a picture), the values of the gadgets will immediately be set up to the new values.

Number of Repeats

This integer gadget controls how often MainActor will play an animation. If it contains for example a "1", MainActor will play it one time. Possible values are 1 to 9999.

This gadget is only accessable if the source project contains an animation.

Flash

If the flash gadget is enabled, MainActor will flash the display while playing animations when the animation format is not fast enough to decompress the current frame within the current timeframe. This means that you can profile your animations using this gadget along with the Frame Info option. The flash is done by shortly changing color 0 of your animation to red. It is only available if the source project contains an animation. Consider:

− Nearly no bigger animation can decompress every picture in 1/60 of a second, dont worry if your animations flashes much (press F1 to get max. speed). If the animations flashes VERY much at 2/60 (press F2) or 3/60 (F3), you should try another (faster) animation format.

− Flashing slowes the animation down in itself, if you have a lot of flashing

it will play much slower. − The Frame Info option also slowes animations down, you get less flash-

ing if you turn it off. − Non-Looping animations will often flash while decompressing picture 1

(IFF-Anim5,IFF-Anim7,IFF-Anim8). − You can easily find out the weak spots of your animations (much flash-

ing), try to smooth the deltas at these points if you use delta compression. (IFF-Anim5,IFF-Anim7,IFF-Anim8).

5.4 System Preferences Window

This window handles everything concerning the outer appearance of Main-Actor.

Select Font

Choose your font for MainActor. This font will then be used for all windows, gadgets, menus. Only the list items of the Section 5.1 [Project Window], page 8 will be displayed in a different font, you can select it in the Section 5.5 [MainActor Preferences Window], page 17.

The actual font and its size are displayed beneath the gadget. The Select Font gadget is not activated when you clone the workbench screen.

Select Display

ONLY AVAILABLE AT OS 2.1 +

Choose your display mode for your MainActor screen. The current display mode is displayed beneath the gadget.

This gadget is only activated when

• MainActor has a screen of its own

• You dont clone the workbench screen

Use own Screen

If enabled, MainActor uses a screen of its own, if not, MainActor will use the workbench screen.

SHANGHAI Windows

If enabled, MainActor uses the SHANGHAI public screen mode, that means that for example shell windows will automatically pop up on the MainActor screen.

This gadget is only activated when you run MainActor on a screen of its own.

Screen Size

This gadget is only available when you have enabled the Use own Screen gadget. You can choose MainActors screen size and its overscan mode with this cycle gadget. The screen size of the current mode is displayed in the two integer gadgets to the right of the Screen Size gadget. The screen size depends on the screen mode you selected with the Select Display gadget.

You can select the following screenmodes: WorkBench

Clones the workbench screen properties (screen size, overscan mode, display mode, font)

Custom Enter MainActors screen size in the two integer gadgets to the right.

Standard Use the standard screen size, without overscan ( for Example 640 * 512) at Pal: HighRes Interlace)

Text Overscan The normal overscan mode which the workbench uses

Video Overcan Tries to open as big a screen as possible. Flickers on normal displays at the right border

Max Overscan Slightly smaller than the "Video Overscan" mode, but as useless on normal displays due to its flicker

Palette Size

Select the colors of your palette with this gadget. The value left to the slider gadget shows the current colors available on your MainActor screen.

This gadget only makes sense if you display pictures or play animations in the window mode. More colors == better dithering of your frames/ pictures. (the window mode is only available under OS 3.0 +).

This gadget is only activated when

• MainActor has a screen of its own

• You dont clone the workbench screen.

Use System Preferences

If you want to enable the changes you have made to your display just click on this gadget.

5.5 MainActor Preferences Window

This window lets you choose/enter some special MainActor attributes.

Select Default Load Module

Lets you choose the default load module. Everytime you open a new project window, the default load module will be the one choosen for this new projects load module.

Select Default Save Module

Lets you choose the default save module. Everytime you open a new project window, the default save module will be the one choosen for this new projects save module.

Select List Item Font

Choose the font for your list items in your Section 5.1 [Project Window], page 8, this font must have a fixed with, the current font and its size are displayed to the right of the gadget, the default is topaz.font 8.

Number of List Items

You can enter how many list items will be displayed in the project windows with this integer gadget, the default is 8.

Save Icon

If you enable this gadet, MainActor will save an workbench-icon for every animation/picture you create. You can then view them by double clicking on their icons. The icons support the same tooltypes as MainView, see Chapter 8 [MainView], page 29. The initial settings of the tooltypes clone the settings of the source project (from which you create the new animation/picture).

Use MainActor Preferences

Will enable the changes made to the MainActor preferences.

5.6 Text Buffer Window

This window shows and buffers messages for feedback of the user. You can use the CURSOR UP/DOWN keys to scroll the text one line up/down. CONTROL+CURSOR UP/DOWN will move to the start/end of the buffer. You can also use the windows scrollbar to scroll through the buffer.

See Section 4.2 [Text Buffer Menu], page 6, for more functions concerning the text uffer.

6 The Modules

MainActor has animation and picture modules of which each has loader and saver modules.

6.1 The Animation Modules

6.1.1 Anim Universal

This module examines the animation format and executes the proper animation module. Pecularities:

The Append function will not work if the destination project has an Universal loader.

6.1.2 Anim Universal PicassoII

Works in the same way as the Universal module but displays all pictures or animations directly on the Picasso-II board from VillageTronic. Pecularities:

It is faster to view animations of the bitmap formats (like Anim-5,7,8) on the native chipset as these formats have to be converted to the picassos chunky pixel mode.

The frame info, display mode etc. options have of course no meaning for this module, the handling of the time codes is however correct and you can still stop the animation and flip to the next frame like normal. The function keys also control the speed of the animation.

6.1.3 IFF-Anim5

The IFF-Anim5 specification was created by Electronic Arts.

It is a "good at compressing but slow at playing" format. Pecularities:

You cannot append to looping animations. You can bypass this by resaving it to a non-looping animation and then appending the frames. You can make a non-looping Animation to a looping animation by appending two frames which are identical to the first two. (You have to reload the animation to make the loop appear) There are some bad animation packers out there, who do not save the first ANHD chunk before the BODY chunk. In this case MainActor cannot insert the timecode for the first frame and will give you an error message (the other timecodes will be set). if you want to time the first frame, just resave the animation with MainActor to avoid this problem.

6.1.4 IFF-Anim7

The IFF-Anim7 specification was created by Wolfgang Hofer.

It is a quit good compromise between speed and animation size. The IFF-Anim7 16 modules use words (16 bits) the Anim-7 32 modules longwords (32 bits) for compression. Therefore the 16 bit variant compresses slightly better than the 32 bit modules, but playback is slower.

The IFF-Anim7 32 modules are, in my opinion, the best choice for timeable, fast animations. Pecularities:

You cannot append to looping animations. You can bypass this by resaving it to a non-looping animation and then appending the frames. You can make a non-looping animation to a looping animation by appending two frames which are identical to the first two. (You have to reload the animation to make the loop appear) There are some bad animation packers out there, who do not save the first ANHD chunk before the BODY chunk. In this case MainActor cannot insert the timecode for the first frame and will give you an error message (the other timecodes will be set). if you want to time the first frame, just resave the animation with MainActor to avoid this problem. The modules do NOT support animations with a DLTA chunk instead of a BODY chunk for the first frame.

6.1.5 IFF-Anim8

The IFF-Anim8 specification was created by Joe Porkka of ASDG. There is no difference in playback speed between the Anim7 and Anim8 formats. Pecularities:

You cannot append to looping animations. You can bypass this by resaving it to a non-looping animation and then appending the frames. You can make a non-looping animation to a looping animation by appending two frames which are identical to the first two. (You have to reload the animation to make the loop appear) There are some bad animation packers out there, who do not save the first ANHD chunk before the BODY chunk. In this case MainActor cannot insert the timecode for the first frame and will give you an error message (the other timecodes will be set). if you want to time the first frame, just resave the animation with MainActor to avoid this problem. The modules do NOT support animations with a DLTA chunk instead of a BODY chunk for the first frame.

6.1.6 IFF-AnimBrush

The IFF-AnimBrush specification was created by Dan Silva of Electronic Arts. Anim Brushes are used/saved by DPaint. Pecularities:

There is no saver module at the moment.

6.2 The Picture Modules

6.2.1 Picture Universal

This module examines the picture format and executes the proper picture module.

6.2.2 Picture Universal PicassoII

Works in the same way as the Universal module but displays all pictures directly on the Picasso-II board from VillageTronic. Pecularities:

The frame info, display mode etc. options have of course no meaning for this module. The only options you have is to press space for the next picture or press escape to abort viewing.

6.2.3 IFF

This is the standard Interchangeable File Format (IFF), the specification

was designed by Electronic Arts.

Pecularities:

Starting from MainActor Version 1.02, the IFF modules can load/save IFF24 (24-Bit) pictures.

6.2.4 SFCP PicassoII

Super Fast Chunky Pixels, the format was created by myself to view pictures more quickly on my Picasso-II board. It was designed for maximal speed on chunky pixel devices. The advantages of this format are :

SFCP uncompresses pictures about 5 to 10 times faster as IFF, at 24 bit it is blazingly fast. SFCP compresses pictures to a size about 10 to 20 percent smaller than IFF in general.

The disadvantages are :

You need a chunky pixel device for fast viewing. It is currently only working on the Picasso-II board. SFCP compression is not as good as IFF on 24 bit pictures, but speedup is about 10 times faster. Pecularities:

SFCP pictures will currently only be recognized by the Universal_ PicassoII universal module as there is currently no support for viewing SFCP pictures on the native chipset. Convert the SFCP pictures you want to view to the IFF format with MainActor if you dont have a supported gfx board.

6.2.5 PCX

PCX is a picture format heavily used on pc’s, it was designed by ZSoft. Pecularities:

No save module yet

6.2.6 Workbench-Icon

These modules support the workbench icons. Pecularities:

The saver module will append ".info" to the save name.

7 The Arexx Interface

MainActors arexx port name is MAINACTOR. Format: <> Parmaters in these angle brackets have to be supplied [] Parameters in square brackets are optional | Vertical bars mean OR and are used to seperate options of which

you can choose one : [ON | OFF] The result variable contains the result of the operation, its possible con-

tents are listed in the command reference.

Look at the MainActor:rexx directory for examples. Note that MainActor does not support blanks in filenames which are arexx arguments. The arexx commands integrated in MainActor 1.16 are:

7.1 Append

Format: Append WILL NOT BE USEABLE WITHOUT A KEYFILE

Appends the selected frames/pictures of the source project to the destination project. result: -

7.2 DeSelectAll

Format: DeSelectAll

Deselects all frames/pictures in the source project. result: -

7.3 DeSelectRange

Format: DeSelectRange <from> <to>

Will deselect all frames/pictures in the source project from the integer value from to to result: -

7.4 DeSelectRangeRequest

Format: DeSelectRangeRequest

The DeSelectRange requester will appear,letting you choose the frames or pictures to deselect. result: -

7.5 GetSPInfo

Format: GetSPInfo Returns information about the current source project: The width of the source project, followed by its height, followed by the

amount of Colors (HAM6,HAM8,EHB will be given back literally for AdPro compatibility), followed by the number of frames/pictures in your project, followed by CACHING=NO or CACHING=YES depending if you have caching en/disabled, followed by LOOP=YES or LOOP=NO depending on the loop status and at last CPF=YES or CPF=NO depending on your color per frame status.

Picture lists have LOOP disabled and CPF enabled. result: width height colors #ofpics caching loop cpf

Examples:

result=320 256 HAM6 220 CACHING=YES LOOP=YES CPF=NO result=320 256 16 10 CACHING=YES LOOP=YES CPF=YES

7.6 GetSPLoader

Format: GetSPLoader

Returns two arguments, the first is ANIM or PIC depending of the loader module type of the source project. The second argument is the name of the module itself. result: ModuleType ModuleName

7.7 GetSPName

Format: GetSPName

If the source project has an animation loaded, it returns the full animation name with path, else it returns two arguments, the first is the name of the first picture with full path, the second is the name of the last picture without its path. result: Name of animation or name of first and last picture.

7.8 GetSPSaver

Format: GetSPSaver

Returns two arguments, the first is ANIM or PIC depending of the saver module type of the source project. The second argument is the name of the module itself. result: ModuleType ModuleName

7.9 LoadProject

Format: LoadProject <Name1> [Name2]

LoadProject will load an animation or picture list, according to the type of the load module of the source project.

Name1 is the name of the animation or the name of the first picture of the picture list.

Name2 is only needed when you want to load a picture list, it is the name of the last picture.

Examples:

LoadProject Graphics:CountAnim.00001 Graphics:CountAnim.00010 LoadProject Graphics:CountAnim.00001 CountAnim.00010 LoadProject Graphics:Test/mandel7 32

result:-

7.10 OpenNewProject

Format: OpenNewProject [CACHING]

Opens a new project window (up to five) If you supply the CACHING parameter, the project window will have caching enabled, otherwise it will be disabled. result: -

7.11 PlayProject

Format: PlayProject

Plays the animation contained in the source project, will fail if the source project is a picture list. result: -

7.12 Quit

Format: Quit

Closes all projects and quits MainActor. result: -

7.13 PrintTxt

Format: PrintTxt <text>

Displays text in the Section 5.6 [Text Buffer Window], page 17. result: -

7.14 PrintAndStoreTxt

Format: PrintAndStoreTxt <text>

Displays and buffers text in the Section 5.6 [Text Buffer Window], page 17. result: -

7.15 RequestInteger

Format: RequestInteger <InitValue> <text>

This command pops up an integer requester, text will be displayed above the integer value which was initialized with InitValue.

Example:

RequestInteger 640 "Enter new Width" result: integer value

7.16 Save

Format: Save <Name> [CPF=YES|NO] [LOOP=YES|NO] WILL NOT BE USEABLE WITHOUT A KEYFILE An animation saver will save the selected frames into a new animation

under the name you put in the Name parameter. CPF decides wether you want a color palette per frame or not, LOOP decides if you want to have a looping animation. LOOP and CPF overrule the normal behaviour, if you dont supply them, everything will be saved as though you had activated the Save gadget.

If you have a picture saver, Name is the base name for the pictures, LOOP and CPF have, in this case, no meaning. result: -

7.17 ScreenToBack

Format: ScreenToBack

Flips the MainActor screen to the back. result: -

7.18 ScreenToFront

Format: ScreenToFront

Flips the MainActor screen to the front. result: -

7.19 SelectAll

Format: SelectAll

Selects all frames/pictures in the source project. result: -

7.20 SelectRange

Format: SelectRange

Will select all frames/pictures in the source project from the integer value

from to to

result: -

7.21 SelectRangeRequest

Format: SelectRangeRequest

The SelectRange requester will appear, letting you choose the frames or pictures to select. result: -

7.22 ShowPictures

Format: ShowPictures

Shows the selected pictures/frames of the source project. result: -

7.23 SetSPLoader

Format: SetSPLoader <ModuleType> <ModuleName>

Sets the load module of the source project to the module type ModuleType with the module name ModuleName.

EXAMPLES:

SetSPLoader PIC IFF

SetSPLoader ANIM IFF-Anim5 result: -

7.24 SetSPSaver

Format: SetSPSaver <ModuleType> <ModuleName>

Sets the save module of the source project to the module type ModuleType with the module name ModuleName.

EXAMPLES:

SetSPSaver PIC IFF

SetSPSaver ANIM IFF-Anim5 result: -

7.25 SetTimecode

Format: SetTimecode <Timecode>

Sets Timecode (a value from 1 to 999) to the selected frames of the source project. result: -

7.26 SwapDeActivatedToSource

Format: SwapDeActivatedToSource

Flips the first deactivated project to the source and the source project to the destination project. result: -

7.27 SwapProjects

Format: SwapProjects

Flips the destination project to the source and the source project to the destination project. result: -

7.28 UnloadProject

Format: UnloadProject

Unloads the current project of the source project. result: -

7.29 Version

Format: Version

Will return the actual version string of MainActor. result: MainActor xx.xx (xx.xx.xx)

7.30 WBToFront

Format: WBToFront

Flips the WorkBench screen to the front. result: -

8 MainView

MainView is MainActors external player. It can display all animations/pictures for which MainActor has the proper loader modules.

The key commands while playing animations or showing pictures are the same as if you would view them directly with MainActor. Thats the reason for not listing them here again, have a look at the Section 5.2 [Project Control Panel Window], page 10 under Play and Show.

If MainActor saved an icon for the animations/pictures to view, Main-View will take the view settings from this icon (overruling its own settings) ! You can change this behaviour through the IgnoreIcons option.

You can start MainView from the workbench or form the cli.

Starting MainView from the Workbench

If you start MainView from the wb (per icon) you can select the file(s) to view in an asl-requester. For multi selecting just press one of the shift keys and select the files with the mouse. (Starting from OS 2.1+) MainView supports the following tooltypes:

Window This tooltype enables the window mode, all animations or pictures will be displayed in an window on the current public screen. It is disabled by default.

FrameInfo FrameInfo enables the FrameInfo mode. It Is disabled by de-

fault.

NoCaching Disables caching, with caching disabled MainView plays direct from the harddisk (or whatever). This allows you to play big animations even if you have not much ram. It is disabled by default.

IgnoreIcons Ignores the icons of the animations/pictures It is disabled by default.

Flash Enables the Flash option. It is disabled by default. XOffset=<Value>

You can directly set the start x offset. Defaults to 0.

YOffset=<Value>

The same for the y offset. Defaults to 0.

Repeat=<Value> Repeats the animation <Value> times. For example Repeat=1 would play the animation one time. Defaults to 100.

Directory=<Path> This is the initial directory for the asl-requester.

Anim Module=<Module> Chooses the default loader module for animations. Default is the Universal module.

Pic Module=<Module> Chooses the default loader module for pictures. Default is the Universal module. If you want to view all pictures for example on the PicassoII gfx board, choose the Universal_PicassoII module.

Starting MainView from the cli

Starting MainView from the cli will give you enhanced functionality. You can for example define patterns for animations or picture lists.

Usage is : MainView [-<options>] <file>

You can enter as many files as you like. MainView supports these options :

    -w           : enables the window mode
    -fi          : enables the frameinfo option
    -nc          : turns caching off
    -i           : ignore the icons for the animations/pictures
    -f           : enables the flash option
    -xvalue      : value describes the initial x offset
    -yvalue      : the same for the y offset.
    -rvalue      : value describes how often the animation
                   should be repeated
    -amodule     : where module is the default animation loader
                   for example -aUniversal selects the universal
                   loader (default).
    -pmodule     : where module is the default picture loader
                   for example -aUniversal selects the universal
                   loader (default).
                   If you want to remap your pictures to the
                   PicassoII board enter -pUniversal_PicassoII

Examples:

MainView -nc -r1 Graphics:Pics/#? Animation/#?8_32 will show all files

in the Graphics:Pics/ directory as well as all files in the Animation/ directory which end with ’8 32’ and all animations will be directly played from the harddisk (one time).

MainView -pUniversal_PicassoII Graphics:#?_24Bit will display all pic-

tures which names end with ’ 24Bit’ on the Picasso-II board.

9 Tutorial

This tutorial will try to (quickly) explain the tasks you can perform with MainActor. It is just meant to give you the idea behind MainActor, you have still to read all of the documentation to get the most out of MainActor.

ALL SAVE OR APPEND FUNCTIONS ARE DISABLED IF YOU DONT HAVE A KEYFILE (DONT HAVE REGISTERED MAINACTOR YET). THE TASKS YOU CANNOT PERFORM IF YOU DONT HAVE A KEYFILE ARE MARKED WITH AN *.

This tutorial will tell you how to

9.1 Load Animations

You need an unused Project to load an animation into MainActor. Select the Select Load Module gadget of the Section 5.1 [Project Window], page 8 and select the proper animation format (if you, for example, want to load an Section 6.1.3 [IFF-Anim5], page 18 animation, select the IFF-Anim5 loader or the Universal loader). If you want to cache the animation (buffer the animation in memory), enable the Caching gadget. Now select the Load gadget of the Section 5.2 [Project Control Panel Window], page 10 and select the animation in the opening asl requester. MainActor will now examine the animation and load it into memory (if caching is enabled). After that the frames of the animation are listed in the scroll list area of the project window, also all gadgets (except the Load and perhaps the Append gadget) of the Project Control Panel Window are now enabled.

9.2 Load Picture(s)

You need an unused Project to load a picture or a picture list into Main-Actor. Select the Select Load Module gadget of the projects window and select the proper picture format (if you, for example, want to load an Section 6.2.3 [IFF], page 20 picture, select the IFF loader). If you want to cache the picture(s) (buffer the picture(s) in memory), enable the Caching gadget. Now select the Load gadget of the Section 5.2 [Project Control Panel Window], page 10 and select the first picture of the picture list in the opening asl requester, if you want to load only one picture, you can cancel the second requester, otherwise select the last picture of the picture list in the second requester. MainActor will now load the selected picture(s) and buffer them into memory (if caching is enabled). After that the picture(s) are listed in the scroll list area of the project window, also all gadgets (except the Load and Play and perhaps the Append gadgets) of the Project Control Panel Window are now enabled.

9.3 Play Animations

You have to load an animation first. Just click on the Play gadget of the Section 5.2 [Project Control Panel Window], page 10. For further details have a look into the documentation.

9.4 Show Pictures

You have to load an animation or a picture list first. Select the pictures you want to show in the source projects list view gadget, after that simply press the Show gadget of the Section 5.2 [Project Control Panel Window], page 10. For further details have a look into the documentation.

9.5 Create/Convert Animations

You have to load an animation or a picture list first. Select the pictures who shall be included in the new animation in the source projects list view gadget. After that select the Select Save Module gadget and select the proper animation format. Now click on the Save gadget of the Section 5.2 [Project Control Panel Window], page 10 and enter the name of the new animation. MainActor will then create the animation, you may have to answer some requesters (looping or non looping animation and so on). For further details have a look into the documentation.

9.6 Save Pictures From Animations

The procedure is exactly the same as before, just select a picture save module instead of an animation module. For further details have a look into the documentation.

9.7 Convert Picture Formats

The procedure is exactly the same as before, just load a picture or a picture list and select another picture formats save module.

9.8 Append To Animations

Load an animation or picture list into the source project and an animation into the destination project. The name of the source projects saver module has to equal the name of the destination projects saver module (Beware of the Universal loader !!) The Append gadget of the Section 5.2 [Project Control Panel Window], page 10 should now be enabled. Select the pictures you want to append (to the destination project) in the source project. After that select the Append gadget, MainActor will now append the selected pictures of the source project to the animation in the destination project. For further details have a look into the documentation.

9.9 Time Animations

Load an animation and select the pictures you want to give a new timecode. Now select the Select Timecode item in the Section 4.3 [Miscellaneous Menu], page 6 and enter the new timecode into the integer gadget. MainActor will now write the new timecodes to the animation, you can check the new timecodes by playing the animation.

9.10 Execute Arexx Scripts

You can directly execute arexx scripts by selecting the Start Arexx Script item in the Section 4.4 [Arexx Menu], page 7 and selecting the arexx scripts. You can map arexx scripts to the function keys through the Install Arexx Scripts item (again in the Arexx Menu). You can really speed up your work by using them, take a look at their header to get a short description what they do.

9.11 Show Pictures/Animations On The PicassoII

Just use the Universal_PicassoII loader modules, they remap the output of every module to the picasso board. Note that it is faster to view animations of the bitmap formats (like Anim-5,7,8) on the native chipset as these formats have to be converted to the picassos chunky pixel mode.

10 Registration

The registration fee is $50 (85 german marks) or any aquivalent sum in your local currency.

You will get your keyfile per disk or email. If you have an email address, please include it in your letter.

When the printed manuals are ready, you will get one. If you dont have ftp access and cant get the updates this way, two free (disk based) updates are included in the shareware fee. You have to contact me everytime you want me to send you an upgrade. You can send cash or international money order to

Markus Moenig
Bruesseler Ring 61
52074 Aachen
Germany
Tel.: (49)-241-71844

If you prefer direct transfer:

Markus Moenig
NO. 20011672
Stadtsparkasse Aachen
BLZ 390 500 00 (Germany)
Usage: MainActor
PLEASE INCLUDE YOUR NAME ON THE TRANSFER ORDER !!

If you have questions/suggestions feel free to contact me through email:

Fido        2:242/7.13
Internet    moenig@pool.informatik.rwth-aachen.de
ADSP        markusm@tolkien.adsp.sub.org

11 The Future

Future enhancements which are on my list: − an animation loader/saver for VRLIs Vista Pro format (VANIM) − mpeg and Jpeg loader − more picture and animation modules (in general) − user defineable shortcuts ala TurboText − support of more 24 Bit modules, for example more modules for graph-

iccards − more powerful arexx commands (for example : syncing arexx commands

to animations) − far more arexx scripts (sinus (de)-acceleration scripts and more) − German documentation − sending printed manuals to registered users − the whole MultiMedia stuff ...

In the long run i want to make a complete MultiMedia package out of MainActor with Sound and Effect modules. The planned sample per frame option is the first step in that direction, instead of integrating the samples in the animations/pictures itself, there will be a new module type : Sound. You will be able to associate sound modules to every frame/picture. Sound modules could for example be samples or protracker play routines.

12 Credits

This package was developed by usage of the following products: − Dice, from Obvious Implementation Corp − Devpac 3.04, from HiSoft − WShell2.0 and Arexx, from Wishful Thinking Corp − TurboText from Oxxi − ToolManager from Stefan Becker − MultiTool II from Boris Jakubaschk Thanks for producing high quality software. Special thanks to: − Stefan Becker for his Amiga knowledge, without him this project would

not have been become possible − all GAMMA and BETA testers − Norbert Bogenrieder for the icons − Paul van der Heu for the original Amigaguide conversion of the docs − Jesper Juul for his optimizing suggestions Index 37

Index

A

Amiga . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Anim Universal . . . . . . . . . . . . . . . . . . . . . . . 18 Anim Universal PicasssoII. . . . . . . . . . . . . 18 Append . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 Append To Animations . . . . . . . . . . . . . . . 32 Arexx Menu. . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 ASDG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

C

Convert Picture Formats . . . . . . . . . . . . . . 32 Create/Convert Animations . . . . . . . . . . . 32 Credits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36

D

Dan Silva . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 DeSelectAll . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 DeSelectRange . . . . . . . . . . . . . . . . . . . . . . . . 22 DeSelectRangeRequest . . . . . . . . . . . . . . . . 23

E

Electronic Arts . . . . . . . . . . . . . . . . . . . . 18, 20 Execute Arexx Scripts. . . . . . . . . . . . . . . . . 33

F

Features . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

G

GetSPInfo . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 GetSPLoader . . . . . . . . . . . . . . . . . . . . . . . . . 23 GetSPName. . . . . . . . . . . . . . . . . . . . . . . . . . . 23 GetSPSaver . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

I

IFF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 IFF-Anim5. . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 IFF-Anim7. . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 IFF-Anim8. . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 IFF-AnimBrush . . . . . . . . . . . . . . . . . . . . . . . 20 Installation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

L

Load Animations. . . . . . . . . . . . . . . . . . . . . . 31 Load Picture(s) . . . . . . . . . . . . . . . . . . . . . . . 31 LoadProject . . . . . . . . . . . . . . . . . . . . . . . . . . 24

M

MainActor Preferences Window . . . . . . . 17 MainView. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 Miscellaneous Menu . . . . . . . . . . . . . . . . . . . . 6

N

New features in v1.1. . . . . . . . . . . . . . . . . . . . 4

O

OpenNewProject . . . . . . . . . . . . . . . . . . . . . . 24 OS 3.0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

P

PCX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 Picture Universal. . . . . . . . . . . . . . . . . . . . . . 20 Picture Universal PicassoII . . . . . . . . . . . . 20 Play Animations . . . . . . . . . . . . . . . . . . . . . . 32 PlayProject . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 PrintAndStoreTxt. . . . . . . . . . . . . . . . . . . . . 25 PrintTxt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 Project Control Panel Window . . . . . . . . 10 Project Menu . . . . . . . . . . . . . . . . . . . . . . . . . . 5 Project View Settings Window . . . . . . . . 13 Project Window. . . . . . . . . . . . . . . . . . . . . . . . 8

Q

Quit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

R

Registration. . . . . . . . . . . . . . . . . . . . . . . . . . . 34 RequestInteger . . . . . . . . . . . . . . . . . . . . . . . . 25

S

Save . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 Save Pictures From Animations . . . . . . . 32 ScreenToBack . . . . . . . . . . . . . . . . . . . . . . . . . 26

Index 38

ScreenToFront . . . . . . . . . . . . . . . . . . . . . . . . 26 SelectAll . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 SelectRange . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 SelectRangeRequest . . . . . . . . . . . . . . . . . . . 26 SetSPLoader . . . . . . . . . . . . . . . . . . . . . . . . . . 27 SetSPSaver . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 SetTimecode . . . . . . . . . . . . . . . . . . . . . . . . . . 27 SFCP PicassoII . . . . . . . . . . . . . . . . . . . . . . . 20 Show Pictures. . . . . . . . . . . . . . . . . . . . . . . . . 32 Show Pictures/Animations On The

PicassoII . . . . . . . . . . . . . . . . . . . . . . . . . 33 ShowPictures . . . . . . . . . . . . . . . . . . . . . . . . . 26 Super Fast Chunky Pixels . . . . . . . . . . . . . 20 SwapDeActivatedToSource . . . . . . . . . . . . 27 SwapProjects . . . . . . . . . . . . . . . . . . . . . . . . . 27 System Preferences Window . . . . . . . . . . . 15

T

Text Buffer Menu . . . . . . . . . . . . . . . . . . . . . . 6 Text Buffer Window . . . . . . . . . . . . . . . . . . 17 The Animation Modules. . . . . . . . . . . . . . . 18 The Arexx Interface . . . . . . . . . . . . . . . . . . . 22 The Future . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 The Menus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 The Modules . . . . . . . . . . . . . . . . . . . . . . . . . . 18 The Picture Modules . . . . . . . . . . . . . . . . . . 20 The Windows . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Time Animations . . . . . . . . . . . . . . . . . . . . . 33 Tutorial. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

U

UnloadProject . . . . . . . . . . . . . . . . . . . . . . . . 28

V

Version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

W

Warm Up . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 WBToFront . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 What is MainActor ? . . . . . . . . . . . . . . . . . . . 3 Windows Menu . . . . . . . . . . . . . . . . . . . . . . . . 7 Wolfgang Hofer . . . . . . . . . . . . . . . . . . . . . . . 19 Workbench-Icon. . . . . . . . . . . . . . . . . . . . . . . 21

Z

ZSoft . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

Table of Contents

1 Warm Up. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Installation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 3 Introduction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

3.1 What is MainActor ? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3.2 Features . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3.3 New features in v1.1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

4 The Menus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

4.1 Project Menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 4.2 Text Buffer Menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 4.3 Miscellaneous Menu. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 4.4 Arexx Menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 4.5 Windows Menu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

5 The Windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

5.1 Project Window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 5.2 Project Control Panel Window . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 5.3 Project View Settings Window . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 5.4 System Preferences Window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 5.5 MainActor Preferences Window . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 5.6 Text Buffer Window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

6 The Modules. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

6.1 The Animation Modules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 6.1.1 Anim Universal. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 6.1.2 Anim Universal PicassoII . . . . . . . . . . . . . . . . . . . . . . . . 18 6.1.3 IFF-Anim5 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 6.1.4 IFF-Anim7 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 6.1.5 IFF-Anim8 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 6.1.6 IFF-AnimBrush . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

6.2 The Picture Modules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 6.2.1 Picture Universal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 6.2.2 Picture Universal PicassoII . . . . . . . . . . . . . . . . . . . . . . 20 6.2.3 IFF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 6.2.4 SFCP PicassoII. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 6.2.5 PCX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 6.2.6 Workbench-Icon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

7 The Arexx Interface. . . . . . . . . . . . . . . . . . . . . . . . . . . 22

7.1 Append . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 7.2 DeSelectAll . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 7.3 DeSelectRange. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 7.4 DeSelectRangeRequest . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 7.5 GetSPInfo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 7.6 GetSPLoader . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 7.7 GetSPName . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 7.8 GetSPSaver . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 7.9 LoadProject . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 7.10 OpenNewProject . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 7.11 PlayProject . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 7.12 Quit. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 7.13 PrintTxt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 7.14 PrintAndStoreTxt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 7.15 RequestInteger . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 7.16 Save. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 7.17 ScreenToBack . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 7.18 ScreenToFront . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 7.19 SelectAll . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 7.20 SelectRange . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 7.21 SelectRangeRequest . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 7.22 ShowPictures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 7.23 SetSPLoader . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 7.24 SetSPSaver . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 7.25 SetTimecode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 7.26 SwapDeActivatedToSource. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 7.27 SwapProjects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 7.28 UnloadProject . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 7.29 Version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 7.30 WBToFront . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

8 MainView . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29

9 Tutorial . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

9.1 Load Animations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 9.2 Load Picture(s) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 9.3 Play Animations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 9.4 Show Pictures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 9.5 Create/Convert Animations. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 9.6 Save Pictures From Animations . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 9.7 Convert Picture Formats. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 9.8 Append To Animations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 9.9 Time Animations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 9.10 Execute Arexx Scripts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 9.11 Show Pictures/Animations On The PicassoII. . . . . . . . . . . . . 33

10 Registration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 11 The Future . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 12 Credits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37